# The Serial Scaling Hypothesis

### v2

@@ F.1 Language modeling
  • Each $f_{\theta,n}$ takes as input $n+2$ elements $x_{1},\dots,x_{n},x,t$, and produces an output $f_{\theta,n}(x,t|x_{1},\dots,x_{n})$.
@@
Comment. The theorem holds for any family of score-networks for which a single forward pass is in $\mathsf{TC}^{0}$. This includes, for example, Transformers and state-space models (Merrill & Sabharwal, 2023b; Merrill et al., 2024).
Finally, since a diffusion model may solve a problem only with high enough probability, instead of solving it deterministically, we make the following definition: A prefix language modeling problem is solved with constant probability bound if there exists some $\epsilon>0$, such that for each input token sequence $x_{1},\dots,x_{n}$, let $x_{correct}$ be the correct response, then
$p(x_{correct}|x_{1},\dots,x_{n})>p(x^{\prime}|x_{1},\dots,x_{n})+\epsilon,\quad\forall x^{\prime}\neq x_{correct}.$ (2)
F.2 Main theorem
@@
Suppose there exists a $\mathsf{TC}^{0}$ family of score-networks $f_{\theta,0},f_{\theta,1},\dots$, such that for each $n$ and each $x_{1},\dots,x_{n}$, the function $f_{\theta,n}(x,t|x_{1},\dots,x_{n})$ exactly computes the score function of some distribution $\rho_{x_{0}|x_{1},\dots,x_{n}}$ with bounded first moment: $\mathbb{E}_{x_{0}\sim\rho_{x_{0}|x_{1},\dots,x_{n}}}[\|x_{0}\|]\leq 1$.
If this family solves a prefix language modeling problem at the limit of infinite time Score-Matching with Langevin Dynamics (SMLD) (Song & Ermon, 2019) with constant probability bound, then the problem is in the $\mathsf{TC}^{0}$ class.
Before we begin the proof, we give the intuitive sketch of the proof. It has two parts. The first part uses the fast convergence rate of the Euler–Maruyama method for solving the diffusion equation to show that we do not need many steps to correctly sample from the diffusion model with high probability.
The second part uses a standard derandomization argument to remove the “high probability” qualifier and obtain a deterministic solution. The idea is, for each length $n$, we replicate the network $k(n)$ times. Each network must take one seed. For each choice of $k(n)$ seeds $s_{1},\dots,s_{k(n)}$, we have a particular deterministic model:
@@
Sample $k(n)$ random seeds, and fix them. This provides a deterministic model. Now, we try this deterministic model on every single possible input of length $n$. There are only $\text{vocab}^{n}$ of them. If we set $k(n)\geq\frac{\ln(\#\text{vocab})}{2(p-1/2)^{2}}$, then by the union bound, the probability that the majority vote is correct on all inputs of length $n$ is nonzero. Thus, there exists a specific choice of random seeds $(s_{1},\dots,s_{k(n)})$ that makes the compound model correct on all inputs of length $n$.
@@
This construction is nonuniform precisely in the part where we have only shown the choice of seeds exists. To actually find these seeds may take exponential time.
Proof.
The idea of the proof is simple. We first construct a universal $\mathcal{O}(1)$-bound on how many steps are sufficient for sampling the SMLD within a constant probability bound, then we derandomize it according to (Hajnal et al., 1993), while still remaining within the $\mathsf{TC}^{0}$ class.
Since the score-network exactly computes the score function for $\rho_{0}$, the score-matching error is exactly zero.
By (Li & Yan, 2024), there exists a constant $c>0$ that does not depend on any parameter in the statement of the theorem (a "universal constant"), such that there exists a noise schedule for DDPM that takes $T$ steps of the following kind. Using the schedule and the score-matching function $f_{\theta}$, we sample from a distribution $\rho_{DDPM,T}$. It satisfies the inequality
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq c\frac{d(\log T)^{3}}{T}$ (3)
Now, since $c$ is a universal constant, we can make up another sufficiently large universal constant $c^{\prime}$, and set $T$ to be $cc^{\prime}d^{2}$, to obtain an upper bound
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq\frac{(\log(cc^{\prime}d^{2}))^{3}}{c^{\prime}d}$ (4)
The key is that $T$ increases with $d$, but does not increase with $n$.
Since the growth of $\log^{3}$ is dominated by linear growth, for any $\epsilon^{\prime}>0$, there exists a large enough universal constant $c^{\prime}$, such that
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq\epsilon^{\prime}$ (5)
for all $d=1,2,3,\dots$.
Let $\epsilon$ be the constant probability bound, then setting $\epsilon^{\prime}=\epsilon/2$, we find that $\rho_{DDPM,T}$ already solves the problem with constant probability bound.
Now we can derandomize this family, obtaining a $\mathsf{TC}^{0}$ family of Boolean circuits that solves the problem deterministically. The details of the derandomization method appear in (Hajnal et al., 1993, Prp. 4.2). The big picture is as follows: we remove the probability by hard-coding a magic constant [^6] This is the part where we assume non-uniformity. In complexity theory, this is an “advice string”. per member of the family, such that sampling for polynomially many times, and taking the majority output, would always give the correct output. By Hoeffding’s inequality, such magic constants exist for a large enough polynomial.
Comment. The requirement for exact score-matching is necessary for the following two reasons:
First, the full form of the inequality from (Li & Yan, 2024) is
@@
$TV(\rho_{DDPM,T},\rho_{\hat{x},0})\leq c\frac{d(\log T)^{3}}{T}+c\epsilon_{\text{score}}\sqrt{\log T},$ (6)
where the term $\epsilon_{\text{score}}$ denotes the score-matching error between the true score function of $\rho_{\hat{x},0}$ and the approximation $f_{\theta}$. As this extra term increases with $T$, the proof above does not apply.
Second, if we have no requirement on score-matching, then there is essentially no constraint on the computational power of SMLD, since any Turing machine can be constructed in SMLD with some (extremely non-score-matching) network. The explicit construction is published in (Liu, 2025).
Practically relevant score-networks are intermediate between two extreme cases. We believe that if $f_{\theta}$ is a good enough, but not perfect, score-matching network, then a generalized version of the above theorem still applies. However, finding the right way to quantify the goodness, as well as proving such a generalization, is left as future work.
@@ G.2 Non-parallelizability Results
We quote the following result from Kirousis and Spirakis (Kirousis & Spirakis, 1988).
@@
Although our analysis uses the DO1 problem in particular, there are some other $\mathsf{P}$-complete problems, such as linear programming, that are $\mathsf{P}$-complete to approximate. See for instance (Sahni & Gonzalez, 1976; Serna, 1991; Díaz et al., 1997; Greenlaw et al., 1995) for some examples. This leads us to conjecture that inherent seriality may be a fairly common phenomenon in RL, not particular to the DO1 setup.

### v3

@@ F.4 Main theorem
  • Each $f_{\theta,n}$ takes as input $n+1$ elements $x_{1},\dots,x_{n},t$, and produces an output $f_{\theta,n}(x_{k_{n}+1},x_{k_{n}+2},\dots,x_{n}|t,x_{1},\dots,x_{k_{n}})$.
@@
Comment. The holds for any family of score-networks for which a single forward pass is in $\mathsf{TC}^{0}$. This includes, for example, Transformers and state-space models (Merrill and Sabharwal, 2023b; Merrill et al., 2024).
For notational convenience, we will thenceforth assume that in our task, there is one true solution per problem. If there are multiple true solutions per problem, then our proof shows that the task of “finding at least one true solution for the original task” is $\mathsf{TC}^{0}$.
Since a diffusion model may solve a problem only with high probability, instead of solving it deterministically, we make the following definition:
A task is solved with constant probability bound if there exists some $\delta>0$, such that for each input token sequence $x_{1},\dots,x_{n}$, let $x_{\text{correct}}$ be the correct discrete solution, then
$p(x_{\text{correct}}|x_{1},\dots,x_{n})>p(x^{\prime}|x_{1},\dots,x_{n})+\delta,\quad\forall x^{\prime}\neq x_{\text{correct}}.$ (2)
As noted before, there are 2 places where a diffusion model can incur error in its solution. The first place is due to using discrete time-steps during the backward diffusion process. The second place is due to score-matching error compared to the true forward diffusion process from the solution manifold.
We are ready to state the theorem. We apologize for the amount of complicated epsilon-delta formality in the statement, but it is what it takes to make it rigorous.
@@
Given a task, and a $\mathsf{TC}^{0}$ family of score-networks $f_{\theta,0},f_{\theta,1},\dots$, for solving the task, by SMLD with $f_{\theta,n}(x_{k_{n}+1},x_{k_{n}+2},\dots,x_{n}|t,x_{1},\dots,x_{k_{n}})$.
Let the $n$-th solution manifold of the task be $X_{n}$. Assume:
  1. 1.
    There exists a constant integer $T>0$, a constant real number $\epsilon_{\text{score}}>0$, a small constant real number $\delta>0$, and a sequence of probability distributions $\rho_{0,n}$ on $X_{n}$, such that
  2. 2.
    Letting $c,c_{\epsilon},c_{M}$ be the constants used in the statement of Theorem F.1, $\epsilon=T^{-c_{\epsilon}}$, and $\epsilon_{\text{score},n}$ be the average score modeling error of using $f_{\theta,n}$ to score-match the forward diffusion process on $\rho_{0,n}$,
  3. 3.
    For all $n=1,2,3,\dots$, we have
    $\mathbb{E}_{x_{0}\sim\rho_{0,n}}[\|x_{0}\|_{2}]$ $\leq T^{c_{M}},$
    $\epsilon_{\text{score},n}$ $\leq\epsilon_{\text{score}},$
    $c\left(\frac{d_{\epsilon}(X_{n})}{T}(\log T)^{3}+\epsilon_{\text{score}}\sqrt{\log T}\right)$ $\leq\frac{1}{2}-\delta.$
Then the task is in the non-uniform $\mathsf{TC}^{0}$ class.
More generally, if the family of score-networks is a $\mathsf{TC}^{k}$ family, and assume:
  1. 1.
    There exists a non-negative integer $k$, a sequence $T_{1},T_{2},\dots$ such that $T_{n}=\mathcal{O}((\log n)^{k})$, a sequence $\delta_{n}$ such that $1/\delta_{n}=\mathsf{poly}(n)$, such that
  2. 2.
    Letting $c,c_{\epsilon},c_{M}$ be the constants used in the statement of Theorem F.1, $\epsilon_{n}=T_{n}^{-c_{\epsilon}}$, and $\epsilon_{\text{score},n}$ be the average score modeling error of using $f_{\theta,n}$ to score-match the forward diffusion process on $\rho_{0,n}$,
  3. 3.
    For all $n=1,2,3,\dots$, we have
    $\mathbb{E}_{x_{0}\sim\rho_{0,n}}[\|x_{0}\|_{2}]$ $\leq T_{n}^{c_{M}},$
    $c\left(\frac{d_{\epsilon_{n}}(X_{n})}{T_{n}}(\log T_{n})^{3}+\epsilon_{\text{score},n}\sqrt{\log T_{n}}\right)$ $\leq\frac{1}{2}-\delta_{n}.$
then the task is in the non-uniform $\mathsf{TC}^{k}$ class.
Proof.
The proof of the case for $\mathsf{TC}^{k}$ is essentially the same as the case for $\mathsf{TC}^{0}$, with more cumbersome notations. Thus, we only prove the case for $\mathsf{TC}^{0}$ explicitly.
Using the big list of assumptions, we can apply Theorem F.1 directly, and conclude that, if we use SMLD with $f_{\theta,n}$ as the score network, for $T$ steps, we would sample from a distribution $\rho_{SMLD,0,n}$ that satisfies
$TV(\rho_{0,n},\rho_{SMLD,0,n})\leq 1/2-\delta$
In particular, this means that after discretization, we obtain the correct solution with probability at least $1/2+\delta$. This then implies that the task is solved with constant probability bound. Now we can derandomize this family, obtaining a $\mathsf{TC}^{0}$ family of Boolean circuits that solves the problem deterministically. The details of the derandomization method appear in (Hajnal et al., 1993, Prp. 4.2). It goes as follows:
for each length $n$, we replicate the network $k(n)$ times. Each network must take one seed. For each choice of $k(n)$ seeds $s_{1},\dots,s_{k(n)}$, we have a particular deterministic model:
@@
Sample $k(n)$ random seeds, and fix them. This provides a deterministic model. Now, we try this deterministic model on every single possible input of length $n$. There are only $\text{vocab}^{n}$ of them. If we set $k(n)\geq\frac{\log(\#\text{vocab})}{2(p-1/2)^{2}}$, then by the union bound, the probability that the majority vote is correct on all inputs of length $n$ is nonzero. Thus, there exists a specific choice of random seeds $(s_{1},\dots,s_{k(n)})$ that makes the compound model correct on all inputs of length $n$.
@@
This construction is nonuniform precisely in the part where we have only shown the choice of seeds [^6] In complexity theory, this choice of seeds is an “advice string”. exists. To actually find these seeds may take exponential time. ∎
F.5 Interpretation
At the high level, to satisfy the assumptions, one needs to have a good enough approximate score with an error $\epsilon_{\text{score}}$. Then, choose $T$ (or $T_{n}$) such that
@@ F.5 Interpretation
$c\left(\frac{d_{\epsilon}(X_{n})}{T}(\log T)^{3}+\epsilon_{\text{score},n}\sqrt{\log T}\right)\leq 1/2-\delta$
For constant $d_{\epsilon}(X_{n})$, we have the following interpretations:
Perfect diffusion model is non-uniform $\mathsf{TC}^{0}$. Consider the case if we have a perfect score function $\epsilon_{\text{score}}=0$, we have $c\left(\frac{d_{\epsilon}(X_{n})}{T}(\log T)^{3}\right)\leq 1/2-\delta$. We can select $T=O(d_{\epsilon}(X_{n}))=O(1)$, hence, the task belongs to the non-uniform $\mathsf{TC}^{0}$ class. This is the same finding as Liu (2025).
Good diffusion model is still non-uniform $\mathsf{TC}^{0}$. Consider if the score function is not perfect. Here, the the accumulated error grows slowly with $T$, i.e., $\epsilon_{\text{score}}\sqrt{\log T}$. As $T$ increases, the term $\frac{d_{\epsilon}(X_{n})}{T}(\log T)^{3}$ decreases quickly, while the term $\epsilon_{\text{score}}\sqrt{\log T}$ increases only slowly and eventually dominates, setting a floor on the total error. If the error is not too large, we can still find a constant $T$ to obtain a constant error bound.
Similar interpretations can be obtained for non-uniform $\mathsf{TC}^{k}$ if $d_{\epsilon_{n}}(X_{n})=O((\log n)^{k})$.
There are two scenarios that the theorem’s assumptions are violated and the conclusion of the theorem does not apply.
$\epsilon_{\text{score}}$ is too large. In this case the inequality fails and our theorem no longer applies. This corresponds to a diffusion model that is not doing a good job at its intended task of score matching. We expect this to occur when the diffusion model is poorly trained or under-parameterized, or when the network is not being used as a score-matching diffusion model at all. If it is not used for score matching, SMLD can in principle implement arbitrary Turing computations via a highly non–score-matching network.
The intrinsic dimension of the solution space $d_{\epsilon}(X_{n})$ grows faster than $\mathsf{polylog}(n)$. In this case, the convergence rate of the diffusion model will be slow enough that the task may go beyond non-uniform $\mathsf{TC}^{k}$ if $d_{\epsilon_{n}}(X_{n})=O((\log n)^{k})$. However, it doesn’t immediately imply that the task is in $\mathsf{P}\setminus\mathsf{TC}$, nor the diffusion model can solve an inherently serial problem. It only means that diffusion models solve a problem with polynomial number of intrinsic dimensions with polynomial steps.
@@ G.2 Non-parallelizability Results
We quote the following result from Kirousis and Spirakis (Kirousis and Spirakis, 1988).
@@
Although our analysis uses the DO1 problem in particular, there are some other $\mathsf{P}$-complete problems, such as linear programming, that are $\mathsf{P}$-complete to approximate. See for instance (Sahni and Gonzalez, 1976; Serna, 1991; Díaz et al., 1997; Greenlaw et al., 1995) for some examples. This leads us to conjecture that inherent seriality may be a fairly common phenomenon in RL, not particular to the DO1 setup.